A 29.0 cm long spring is hung vertically from a ceiling and stretches to 35.7 cm when a 7.00 kg mass is hung from its free end.


(a)

Find the spring constant (in N/m).

 __N/m


(b)

Find the length of the spring (in cm) if the 7.00 kg weight is replaced with a 165 N weight.

__ cm


1
Expert's answer
2021-09-25T03:55:26-0400

(a)

"Δx = 35.7-29.0 = 6.7 \;cm = 0.067 \;m \\

Mg=kΔx \\

7.0 \times 9.8 = k \times 0.067 \\

k = \frac{68.6}{0.067}=1023.88 \;N/m"

(b)

"Mg = 165 \;N \\

165 = 1023.88 \times Δx \\

Δx = \frac{165}{1023.88} = 0.1611 \; = 16.11 \;cm"